# Canonical differential equations for the elliptic two-loop five-point integral family relevant to $t\bar{t}+$jet production at leading colour

### v1

@@ Abstract
We present differential equations (DEs) in canonical form for a family of two-loop five-point Feynman integrals containing elliptic functions and nested square roots. This is the only family for which canonical DEs were not yet available among those required to compute the two-loop leading-colour amplitude for top-pair production in association with a jet at hadron colliders. We write the DEs in terms of one-forms having (locally) simple poles at all singular points, and highlight the ‘duplet’ structure that generalises the even/odd charge of square roots to nested roots. All transcendental functions in the DEs are expressed in closed form using complete elliptic integrals, while one-forms free of elliptic functions are given in terms of logarithms, including those with the nested roots. In addition to marking a significant step towards next-to-next-to-leading order QCD predictions for an important LHC process, this work represents the first time that a canonical basis of integrals involving elliptic functions has been obtained for a five-particle process.
@@
One of the most powerful frameworks for computing Feynman integrals consists in viewing them as the solution to certain differential equations (DEs) [1, 2, 3, 4, 5, 6]. The key step is to transform the DEs so that they take a canonical form [6], which makes solving them dramatically easier. In the best understood case, the DEs are canonical if the dependence on the regulator of dimensional regularisation is factorised, and the DEs contain solely logarithmic differential one-forms (${\mathrm{d}}\log$’s). In such a case, we have a well developed mathematical technology to handle the solution (see refs. [7, 8, 9, 10] for reviews), and the state of the art has reached integrals with five [11] and six particles [12, 13, 14] at three- and two-loop order, respectively.
However, this is not the general case [15]: starting at two-loop order, Feynman integrals may involve transcendental functions linked to higher-genus geometries [16], such as periods of elliptic curves, whose DEs cannot be expressed in the form above. Such functions are typically linked to the presence of massive virtual particles. Despite significant recent progress in extending the construction of canonical DEs [17, 18, 19, 20, 21, 22, 19, 23, 24], the most advanced applications remain confined to processes with comparatively simpler four-particle kinematics [25, 26, 19, 27, 28, 29]. Advancing our methodology to handle high-multiplicity processes with massive virtual particles is imperative. Indeed, this is one of the main bottlenecks to keep the theoretical uncertainties in line with the experimental ones for a wide range of processes central to the LHC physics programme [30]. Moreover, thanks to their rich kinematics, high-multiplicity processes offer an extraordinary laboratory for uncovering new properties of QFT.
We break through the wall of elliptic integrals with five-point kinematics, and obtain canonical DEs for a family of Feynman integrals relevant for top-pair production in association with a jet at hadron colliders that involves elliptic functions [31, 32, 33]. Our result completes the construction of canonical DEs for all families required to compute the leading-colour two-loop QCD amplitude for this process, marking a key step towards obtaining next-to-next-to-leading order (NNLO) QCD predictions. This achievement is of immediate importance for the Large Hadron Collider (LHC), where top-pair production with a jet is a high-priority process [34, 35, 36, 37]. Its high sensitivity to the top-quark mass [38, 39, 40, 41] and the increasing experimental precision in fact call for the NNLO QCD corrections to be computed. Furthermore, our results advance our theoretical understanding of elliptic Feynman integrals to a new, high-multiplicity regime, and provide precious analytical data to fuel ongoing formal studies and open new avenues of research.
@@ I Setting the stage
where ${\mathrm{d}}$ is the total differential with respect to $\vec{x}$. The connection matrix ${\mathrm{d}}A(\vec{x},\epsilon)$ for the basis $\vec{\mathcal{I}}$ of ref. [32] depends quadratically on $\epsilon$, and contains non-logarithmic one-forms. These non-canonical features are however localised in the two sectors shown in fig. 2, i.e., in the subsets of integrals having the propagators of the graphs in the figure. More concretely, the ‘homogeneous’ blocks—i.e., the blocks on the diagonal of the connection matrix that couple the MIs of the same sector—were cast into ‘$\epsilon\times{\mathrm{d}}\log$’ form for all sectors but the two in fig. 2.
As noticed in ref. [32], the homogeneous block of the sector in fig. 2(b) is $\epsilon$-factorised by rotating $\mathcal{I}_{19}$ and $\mathcal{I}_{20}$ as
@@
where $r_{k}$ denotes a degree-$k$ polynomial in $\vec{x}$. The resulting $\epsilon$-factorised block was not written in terms of ${\mathrm{d}}\log$’s, and it is unclear whether this is generally possible in the presence of nested roots [45]. In section III, we show that this is indeed possible in our case, and discuss the symmetries associated with the nested square roots.
The sector in fig. 2(a) was instead shown to involve elliptic functions associated with the elliptic curve $y^{2}=\mathcal{P}(z)$,
@@
where the roots $e_{i}$ are functions of $\vec{x}$. In particular, the maximal cut of the scalar integral of this sector (${\cal I}_{35}$, with a suitable normalisation) at $\epsilon=0$ is given, depending on the integration cycle, by the periods of the elliptic curve [32]. We choose them as
@@
$\kappa^{2}=e_{32}e_{41}/(e_{31}e_{42})$, with $e_{ij}=e_{i}-e_{j}$, is the elliptic modulus, and the normalisations in eq. 7 were chosen to simplify the $m_{t}^{2}$ expansion. The presence of the elliptic curve is the main obstacle towards achieving canonical DEs, which we overcome in the next section.
Figure 1: Graph of the integral family studied in this letter. Black (red double) lines are massless (massive).
@@ II Construction of a Canonical Basis
There is no consensus yet on what ‘canonical’ means in addition to the factorisation of $\epsilon$ beyond the ${\mathrm{d}}\log$ case [46, 47, 24]. We adopt a local definition: we require that, for each singular point, it is possible to choose the appearing period of the elliptic curve such that all one-forms have at most simple poles. In this section, we will be concerned with achieving the factorisation of $\epsilon$. We will verify that the resulting DEs fulfil this notion of canonical in section III. We denote by $T(\vec{x},\epsilon)$ the transformation matrix from the basis $\vec{\mathcal{I}}$ of ref. [32] to the new, canonical basis $\vec{\mathcal{J}}$, i.e., $\vec{\mathcal{J}}=T\cdot\vec{\mathcal{I}}$. We include in $T$ the transformation in eq. 4.
We construct $T$ in two steps. First, we transform the homogeneous block of the DEs related to the elliptic sector into $\epsilon$-factorised form. Second, we perform a similar rotation to factorise $\epsilon$ in the inhomogeneous blocks that couple other sectors to the ones shown in fig. 2. In both cases, we follow the approach of ref. [19]. An important part of this method is to find a suitable starting basis that simplifies the DEs as much as is possible without introducing transcendental functions. We find the basis of ref. [32] to be a suitable starting point. In particular, the maximal cut of $\mathcal{I}_{35}$ is related to the periods of the elliptic curve (see appendix A of ref. [32] and eq. 18).
@@
by standard methods. The solutions to this equation are the two periods in eq. (7). We then can convert back to the $2\times 2$ system to write $W$ in terms of the periods.
@@
Achieving $\epsilon$-factorisation therefore amounts to integrating out the remaining elements at $\epsilon=0$ [48]. However, as was the case for several examples in ref. [19], not all elements are total derivatives in $m_{t}^{2}$. As a result, we need to introduce a new function, $G_{1}$, defined through its $m_{t}^{2}$-derivative, as $\partial_{m_{t}^{2}}G_{1}=\bigl{(}A^{\text{ss},\epsilon=0}_{m_{t}^{2},\text{hom.,ell.}}\bigr{)}_{2,1}$, which is linear in $\psi_{1}$ and $\psi_{1}^{\prime}$ (the same as in $W^{\text{ss}}$), and rational in $\vec{x}$.
Having worked out the intricate dependence of the resulting transformation on all kinematic variables (see appendix A), we can successively check the $\epsilon$-factorisation of the homogeneous DEs not only in $m_{t}^{2}$, but also in the other variables. We find that, thanks to our careful choice for the solutions of the appearing differential operators, these are automatically $\epsilon$-factorised, so that no further ($m_{t}^{2}$-independent) transformations are required.
@@
Similarly to the first step, these transformations suffice to $\epsilon$-factorise the DEs in all variables, even though they were derived by only considering the derivative w.r.t. $m_{t}^{2}$. Moreover, while from the point of view of the DEs the new functions $G_{i}$ are defined through their partial derivatives, we were also able to express them explicitly as complete elliptic integrals of the third kind by considering different Baikov representations [49, 50] and performing the integrals appearing in them, see appendix A. This result will be useful in view of the numerical evaluation of the solution, which we leave to future work.
@@ III Analytic Structure
where $M_{i}$ are constant rational matrices, and $\omega_{i}(\vec{x})$ are $\mathbb{Q}$-linearly independent differential one-forms. We verified the linear independence either by treating the elliptic functions ($\psi_{1}$, $G_{i}$) as independent variables, or by evaluating them numerically. We expressed the 83 $\omega_{i}(\vec{x})$’s free of elliptic functions as ${\mathrm{d}}\log W$ for some algebraic function $W$. This required constructing 13 new ${\mathrm{d}}\log$’s involving the nested square roots, besides the ones of ref. [32]. We chose the 29 one-forms containing elliptic functions from the entries of the connection matrix, and wrote them in terms of differentials of $\vec{x}$. The definition of the MIs $\vec{\cal J}$ contains overall (nested) square roots and transcendental functions. This induces two gradings on the one-forms.
First, it is well known that one-forms and MIs can be chosen to be either even or odd upon flipping the sign of each square root. While this holds also for $\sqrt{n_{\pm}}$, flipping the interior root—$\sqrt{\Delta_{5}}$—exchanges $\sqrt{n_{+}}$ and $\sqrt{n_{-}}$, leading to a more complicated behaviour. One way to generalise the even/odd parity to a picture that includes nested square roots is to view it as the trivial/sign representation of $\mathbb{Z}_{2}$, the group formed by the operations of flipping each square root. The relevant group for the nested square roots $\sqrt{n_{\pm}}$ is $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$: one $\mathbb{Z}_{2}$ flips the sign of the exterior root, the other of the interior one. [^2] The mathematical framework is that of Galois theory. The Galois group of a polynomial is, roughly speaking, the group of unique permutations of its roots. Square roots are roots of quadratic polynomials whose Galois group is $\mathbb{Z}_{2}$. The nested square roots ±$\pm\sqrt{n_{\pm}}$ are instead roots of $P(z)=z^{4}+az^{2}+b$ (for $a$, $b$ functions of $\vec{x}$), whose Galois group is $D_{4}\simeq\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. See also refs. [51, 52] for similar applications of Galois theory to Feynman integrals. For the exterior root we have the usual even/odd representation. For the interior root, instead, we have a two-dimensional representation of $\mathbb{Z}_{2}$ generated by the Pauli matrix $\sigma_{1}$, acting on duplets. In other words, we arrange all objects into duplets whose elements transform into each other upon flipping $\sqrt{\Delta_{5}}$. Examples of such duplets are $(\sqrt{n_{+}},\sqrt{n_{-}})$, $(\mathcal{J}_{19},\mathcal{J}_{20})$, $(G_{3,+},G_{3,-})$, and $(\omega_{4},\omega_{5})$. This duplet structure constrains the form of the DEs, as
@@
which manifests the grading. We then constrain the polynomials $c_{i}$ by imposing that the singular locus is that of the target one-form. This amounts to solving quartic equations, for which we adopt ad-hoc methods leveraging the sparsity of the system and of the solution. In some cases, it is easier to integrate the one-form in one variable, setting all others to numbers and rationalising one of the roots, and use the result to constrain the ansatz. Interestingly, the polynomial $C\coloneqq n_{+}n_{-}$ does not appear as the argument of a ${\mathrm{d}}\log$. This observation is relevant in view of the programme aimed at determining the logarithmic one-forms from the Landau equations [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68]: $C$ appears in the Landau discriminant, but only factoring it as $C=(a+b\sqrt{\Delta_{5}})(a-b\sqrt{\Delta_{5}})$, for polynomials $a$ and $b$, reveals $n_{\pm}=a\pm b\sqrt{\Delta_{5}}$ as possible arguments of square roots.
The second grading, that we dub elliptic grading, is associated with the elliptic functions. By grading $+1$ the periods and thus the $G_{i}$’s, all MIs and one-forms gain a uniform grade ranging from $-2$ to $+2$, as already observed in other cases. Note that a function having elliptic grade 00 may still contain elliptic functions, as ratios with $\psi_{1}$ in the denominator also appear, e.g. $G_{1}/\psi_{1}$. We find that elliptic one-forms appear also in the entries of the connection matrix coupling MIs of sectors that are not elliptic on the maximal cut. Moreover, $\psi_{1}^{\prime}$ is absent from the DEs, although present in the transformation $T$.
Finally, we verified that the DEs in eq. 15 are canonical according to the local definition given above. They contain transcendental functions, in particular $\psi_{1}$. What matters for the $\epsilon$-factorisation is that $\psi_{1}$ satisfies the DE (12); any linear combination of $\psi_{1}$ and $\psi_{2}$ would work, and this freedom extends to the $G_{i}$’s. However, in the neighbourhood of certain singular points, one solution to eq. 12 is holomorphic while the other has logarithmic singularities; which one depends on the point. We verified that, for each singular point, there is a solution to eq. 12 such that all one-forms have at most simple poles. In practice, we pull the one-forms back to paths parameterised by one variable, and expand the elliptic functions around each singular point via Frobenius’ method [69, 70], choosing the holomorphic solution.

### v2

# Canonical Differential Equations for the Elliptic Two-Loop Five-Point Integral Family Relevant to $t\bar{t}+$Jet Production at Leading Color

@@ Abstract
We complete the construction of canonical differential equations (DEs) for all families of Feynman integrals appearing in the two-loop leading-color QCD amplitude for top-pair production in association with a jet at hadron colliders. This allows us to obtain analytic results for all required Feynman integrals in terms of iterated integrals with closed-form kernels. To achieve this, we study the two-loop five-point integral family for which canonical DEs were previously unavailable due to the appearance of elliptic functions and nested square roots. In addition to marking a significant step towards next-to-next-to-leading-order QCD predictions for a high-priority LHC process, this is the first time that canonical DEs are obtained for Feynman integrals where elliptic functions appear in conjunction with the high algebraic complexity of a process with more than four particles. Our results also reveal a number of new interesting analytic features, such as a ‘duplet’ structure that generalizes the even/odd parity of square roots to nested square roots.
@@
One of the most powerful frameworks for computing Feynman integrals consists in viewing them as the solution to certain differential equations (DEs) [1, 2, 3, 4, 5, 6]. The insight that a canonical form of the DEs [6]—one which dramatically simplifies the solution—exists has made possible increasingly many results in high-energy physics. In the best understood case, canonical means that the dependence on the regulator of dimensional regularization is factorized, and the DEs contain solely logarithmic differential one-forms (${\mathrm{d}}\log$’s). In such a case, we have a well developed mathematical technology to handle the solution (see Refs. [7, 8, 9, 10] for reviews), and the state of the art has reached integrals with five [11] and six particles [12, 13, 14] at three- and two-loop order, respectively.
However, starting at two-loop order, Feynman integrals may involve transcendental functions linked to higher-genus geometries [15, 16], such as periods of elliptic curves, whose DEs cannot be expressed in the form above. Such functions are typically linked to the presence of massive virtual particles. Despite recent progress in extending the construction of canonical DEs [17, 18, 19, 20, 21, 22, 19, 23, 24], the most advanced applications are confined to processes with comparatively simpler four-particle kinematics [25, 26, 19, 27, 28, 29, 30]. Advancing our methodology to handle high-multiplicity processes with massive virtual particles is imperative. This is one of the main bottlenecks to keep the theoretical uncertainties in line with the experimental ones for many processes in the Large Hadron Collider (LHC) program [31]. Moreover, thanks to their rich kinematics, high-multiplicity processes offer an extraordinary laboratory for uncovering new properties of QFT.
We break through the wall of elliptic integrals with five-point kinematics, and obtain canonical DEs for a family of Feynman integrals relevant for top-pair production with a jet at hadron colliders that involves elliptic functions [32, 33, 34]. Our result completes the construction of canonical DEs for all families appearing in the leading-color two-loop QCD amplitude for this process, allowing us to express all required Feynman integrals analytically in terms of iterated integrals [35], and thus marking a key step towards obtaining next-to-next-to-leading-order (NNLO) QCD predictions. This achievement is of immediate importance for the LHC, where top-pair production with a jet is a high-priority process [36, 37, 38, 39]. Its high sensitivity to the top-quark mass [40, 41, 42, 43] and the increasing experimental precision call for the NNLO QCD corrections to be computed. Our results advance the theoretical understanding of elliptic Feynman integrals to a new, high-multiplicity regime, provide analytical data to fuel ongoing formal studies, and open new research avenues.
@@ I Setting the stage
where $\mathrm{d}$} is the total differential with respect to $\vec{x}$. The connection matrix ${\mathrm{d}}A(\vec{x},\epsilon)$ for the basis $\vec{\mathcal{I}}$ of Ref. [33] depends quadratically on $\epsilon$, and contains non-logarithmic one-forms. These non-canonical features are however localized in the two sectors shown in Fig. 2, i.e., in the subsets of integrals having the propagators of the graphs in the figure. More concretely, the ‘homogeneous’ blocks—i.e., the blocks on the diagonal of the connection matrix that couple the MIs of the same sector—were cast into $\epsilon\times{\mathrm{d}}\log$’ form for all sectors but the two in Fig. 2.
As noticed in Ref. [33], the homogeneous block of the sector in Fig. 2(b) is $\epsilon$-factorized by rotating $\mathcal{I}_{19}$ and $\mathcal{I}_{20}$ as
@@
where $r_{k}$ denotes a degree-$k$ polynomial in $\vec{x}$. The resulting $\epsilon$-factorized block was not written in terms of ${\mathrm{d}}\log$’s, and it is unclear whether this is generally possible in the presence of nested roots [47]. We will show that this is indeed possible in our case, and discuss the symmetries associated with the nested square roots. Note that the interior square root, $\sqrt{\Delta_{5}}$ in Eq. 5, can be rationalised, e.g., by using momentum-twistor variables [48], at the cost of increasing the algebraic complexity of the expressions and obscuring the symmetry properties discussed in Section III. We therefore refrain from doing such a change of variables.
The sector in Fig. 2(a) was shown to involve elliptic functions associated with the elliptic curve $y^{2}=\mathcal{P}(z)$,
@@
where the roots $e_{i}$ are functions of $\vec{x}$. In particular, the maximal cut of the scalar integral of this sector (${\cal I}_{35}$, with a suitable normalization) at $\epsilon=0$ is given, depending on the integration cycle, by the periods of the elliptic curve [33]. We choose them as
@@
$\kappa^{2}=e_{32}e_{41}/(e_{31}e_{42})$, with $e_{ij}=e_{i}-e_{j}$, is the elliptic modulus, and the normalizations in Eq. 7 were chosen to simplify the $m_{t}^{2}$ expansion. The presence of the elliptic curve is the main obstacle towards achieving canonical DEs, which we overcome in the next section.
Figure 1: Graph of the integral family studied in this Letter. Thin (thick) lines are massless (massive).
@@ II Construction of a Canonical Basis
There is no consensus yet on what ‘canonical’ means in addition to the factorization of $\epsilon$ beyond the ${\mathrm{d}}\log$ case [49, 50, 24]. We adopt a local definition: we require that, for each singular point, it is possible to choose the appearing period of the elliptic curve such that all one-forms have at most simple poles. In this section, we will be concerned with achieving the factorization of $\epsilon$. We will verify that the resulting DEs fulfil this notion of canonical in the next Section. We denote by $T(\vec{x},\epsilon)$ the transformation matrix from the basis $\vec{\mathcal{I}}$ of Ref. [33] to the new, canonical basis $\vec{\mathcal{J}}$, i.e., $\vec{\mathcal{J}}=T\cdot\vec{\mathcal{I}}$. We include in $T$ the transformation in Eq. 4.
We construct $T$ in two steps. First, we transform the homogeneous block of the DEs related to the elliptic sector into $\epsilon$-factorized form. Second, we perform a similar rotation to factorize $\epsilon$ in the inhomogeneous blocks that couple other sectors to the ones shown in Fig. 2. In both cases, we follow the approach of Ref. [19]. An important part of this method is to find a suitable starting basis that simplifies the DEs as much as is possible without introducing transcendental functions. We find the basis of Ref. [33] to be a suitable starting point. In particular, the maximal cut of $\mathcal{I}_{35}$ is related to the periods of the elliptic curve (see Appendix A of Ref. [33] and Eq. 18).
@@
by standard methods. The solutions to this equation are the two periods in Eq. 7. We then can convert back to the $2\times 2$ system to write $W$ in terms of the periods.
@@
Achieving $\epsilon$-factorization therefore amounts to integrating out the remaining elements at $\epsilon=0$ [51]. However, as was the case for several examples in Ref. [19], not all elements are total derivatives in $m_{t}^{2}$. As a result, we need to introduce a new function, $G_{1}$, defined through its $m_{t}^{2}$-derivative, as $\partial_{m_{t}^{2}}G_{1}=\bigl{(}A^{\text{ss},\epsilon=0}_{m_{t}^{2},\text{hom.,ell.}}\bigr{)}_{2,1}$, which is linear in $\psi_{1}$ and $\psi_{1}^{\prime}$ (the same as in $W^{\text{ss}}$), and rational in $\vec{x}$.
Having worked out the intricate dependence of the resulting transformation on all kinematic variables (see the Appendix), we can successively check the $\epsilon$-factorization of the homogeneous DEs not only in $m_{t}^{2}$, but also in the other variables. We find that, thanks to our careful choice for the solutions of the appearing differential operators, these are automatically $\epsilon$-factorized, so that no further ($m_{t}^{2}$-independent) transformations are required.
@@
Similarly to the first step, these transformations suffice to $\epsilon$-factorize the DEs in all variables, even though they were derived by only considering the derivative w.r.t. $m_{t}^{2}$. Moreover, while from the point of view of the DEs the new functions $G_{i}$ are defined through their partial derivatives, we were also able to express them explicitly as complete elliptic integrals of the third kind by considering different Baikov representations [52, 53] and performing the integrals appearing in them, see the Appendix. This result will be useful in view of the numerical evaluation of the solution, which we leave to future work.
@@ III Analytic Structure
where $M_{i}$ are constant rational matrices, and $\omega_{i}(\vec{x})$ are $\mathbb{Q}$-linearly independent differential one-forms. We verified the linear independence either by treating the elliptic functions ($\psi_{1}$, $G_{i}$) as independent variables—i.e., we view the entries of the connection matrix as rational functions of $\vec{x}$, $\psi_{1}$ and $G_{i}$—or by evaluating them numerically. We expressed the 83 $\omega_{i}(\vec{x})$’s free of elliptic functions as ${\mathrm{d}}\log W$ for some algebraic function $W$. This required constructing 13 new ${\mathrm{d}}\log$’s involving the nested square roots, besides the ones of Ref. [33]. We chose the 29 one-forms containing elliptic functions from the entries of the connection matrix, and wrote them in terms of differentials of $\vec{x}$. The definition of the MIs $\vec{\cal J}$ contains overall (nested) square roots and transcendental functions. This induces two gradings on the one-forms.
[…]
@@
which manifests the grading. We then constrain the polynomials $c_{i}$ by imposing that the singular locus is that of the target one-form. This amounts to solving quartic equations, for which we adopt ad-hoc methods leveraging the sparsity of the system and of the solution. In some cases, it is easier to integrate the one-form in one variable, setting all others to numbers and rationalising one of the roots, and use the result to constrain the ansatz. Interestingly, the polynomial $C\coloneqq n_{+}n_{-}$ does not appear as the argument of a ${\mathrm{d}}\log$. This observation is relevant in view of the program aimed at determining the logarithmic one-forms from the Landau equations [58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71]: $C$ appears in the Landau discriminant, but only factoring it as $C=(a+b\sqrt{\Delta_{5}})(a-b\sqrt{\Delta_{5}})$, for polynomials $a$ and $b$, reveals $n_{\pm}=a\pm b\sqrt{\Delta_{5}}$ as possible arguments of square roots.
The second grading, that we dub elliptic grading, is associated with the elliptic functions. By grading +1+1 the periods and thus the $G_{i}$’s, all MIs and one-forms gain a uniform grade ranging from 2-2 to +2+2, as already observed in other cases. Note that a function having elliptic grade 0 may still contain elliptic functions, as ratios with $\psi_{1}$ in the denominator also appear, e.g. $G_{1}/\psi_{1}$. We stress that the emergence of this grading is here a mere observation, that we use to classify MIs and one-forms; its implications are left for future study. Additionally, we note that elliptic one-forms appear also in the entries of the connection matrix coupling MIs of sectors that are not elliptic on their maximal cut, and that $\psi_{1}^{\prime}$ is absent from the DEs, although present in the transformation $T$.
Finally, we verified that the DEs in Eq. 15 are canonical according to the local definition given above. They contain transcendental functions, in particular $\psi_{1}$. What matters for the $\epsilon$-factorization is that $\psi_{1}$ satisfies the DE (12); any linear combination of $\psi_{1}$ and $\psi_{2}$ would work, and this freedom extends to the $G_{i}$’s. However, in the neighborhood of certain singular points, one solution to Eq. 12 is holomorphic while the other has logarithmic singularities; which one depends on the point. We verified that, for each singular point, there is a solution to Eq. 12 such that all one-forms have at most simple poles. In practice, we pull the one-forms back to paths parameterized by one variable, and expand the elliptic functions around each singular point via Frobenius’ method [72, 73], choosing the holomorphic solution.